OpenStudy (nuccioreggie):

Compare the table and equation. x y 4 12 5 15 6 18 Equation: y = 3x Which representation has the greatest slope? The equation has the greatest slope. The table has the greatest slope. The table and equation have the same slope. Their slopes cannot be determined.

OpenStudy (michele_laino):

the slope of the line \(y=kx\), is given by the value of \(k\), so, what is the slope of the second graph: \(y=3x\)?

OpenStudy (michele_laino):

hint: what is the coefficient which multiplies the variable \(x\)

OpenStudy (nuccioreggie):

i got x/3

OpenStudy (michele_laino):

please follow my reasoning. Please what is the corùefficient which multiplies \(x\) into the formula \(y=3x\)?

OpenStudy (michele_laino):

oops.. coefficient*

OpenStudy (michele_laino):

the number at the left of \(x\), is...?

OpenStudy (nuccioreggie):

ohh its A

OpenStudy (nuccioreggie):

3

OpenStudy (michele_laino):

no, I think that option A, is a wrong option correct! the slope of the second function is \(3\)!

OpenStudy (michele_laino):

now, we have to compute the slope of the first function: here we can note that: \[\begin{gathered} \frac{{12}}{4} = ...? \hfill \\ \frac{{15}}{5} = ...? \hfill \\ \frac{{18}}{6} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please what is \(12/4=...?\)

OpenStudy (nuccioreggie):

3

OpenStudy (michele_laino):

and \(15/5=...?\)

OpenStudy (nuccioreggie):

3

OpenStudy (michele_laino):

finally: \(18/6=...?\)

OpenStudy (nuccioreggie):

3

OpenStudy (michele_laino):

I think you got the slope of the first function, since every time, we have divided the \(y\) value by the corresponding \(x\) value. So, what is the right option?

OpenStudy (nuccioreggie):

c

OpenStudy (michele_laino):

correct!